let x1=1 and Xn+1=squrrt(3xn)for all n€N show that (xn) is bounded ,monotone and converges to 3

1
Expert's answer
2022-02-24T15:47:01-0500

Let "f(x)=\sqrt{3x}" for all "x>0", then we have "x_{n+1}=f(x_n)" for all "n\in\mathbb{N}".

Since "f(x)" is a monotonously increasing function, then any recurrent sequence satisfying "x_{n+1}=f(x_n)" is monotone. Really:

If "x_0\leq x_1" then "f(x_0)\leq f(x_1)", that is, "x_1\leq x_2" and, by induction, "x_n\leq x_{n+1}".

If "x_0\geq x_1" then "f(x_0)\geq f(x_1)", that is, "x_1\geq x_2" and, by induction, "x_n\geq x_{n+1}".

By the condition, "x_0=1", thus, "x_1=\sqrt{3}>1=x_0", therefore, this sequence is monotonously increasing. Note that if for any "n", "x_n< 3" then "x_{n+1}=f(x_n)<f(3)=3", and, by induction, "x_n< 3" for all "n\in\mathbb{N}". Thus, the given sequence is increasing and bounded from above by 3. Therefore, "x_n" is convergent. Let "a\in\mathbb{R}" is limit of this sequence.

Since "a=\lim\limits_{n\to+\infty}x_n" and "f(x)" is continuous everywhere on "(0,+\infty)", we have

"f(a)=\lim\limits_{n\to+\infty}f(x_n)=\lim\limits_{n\to+\infty}x_{n+1}=a"

"\sqrt{3a}=a" implies "a=0" or "a=3". Since "x_n" is increasing, "\lim\limits_{n\to+\infty}x_n\geq x_0=1", therefore, "a\ne 0", and hence, "a=3".